if you have 5 red chickens, 17 white chickens, three black chickens, and one spotted chicken, what percent is white chickens? I figured it would be approximately 74%. But the answer is 65. Can you show me how this is, please?
Answer by rosey51235(30)

hi:)
To find a percent, (in this problem) you would divide the #of white chickens by the total# of chickens. (also I must insert...where on EARTH would one even accquire a spotted chicken??? xD) anyways......
5+17=22, 22+3=25, and 25+1=26. Total chickens=26.
now divide white chickens by total chickens..... 17/26=.6538461
round it and you get 65%....thats how you get 65%.
how did you manage to get 74%, darling? :)
Hope I helped! :)
-Rose